OpenStudy (anonymous):

find the value b in each perfect square trinomial: 3x^2+bx+27 explain please.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

perfect square are those you can write as \[(tx \pm ry)^2\] When you expand a binomial in tthat form you'll get \[t^2x^2 \pm 2txry + r^2y^2\] In your case, \[r^2y^2=27\rightarrow ry=3\sqrt {3}\] \[t^2x^2=3x^2 \rightarrow tx=\sqrt {3} x\] Then \[2txry= 18x \] In your case "b"=2try So, you'll get b=18 Got it?

OpenStudy (anonymous):

i got how u got 3√3 but then no what did u exactly do?

OpenStudy (anonymous):

The exact same, aply a square root to t^2x^2, to obtain tx

OpenStudy (anonymous):

so 3√3 multiply by what?

OpenStudy (anonymous):

by tx, the easiest explanation says \[(ax+c)^2=a^2+2acx+c^2\], in that case, your "b" will be 2ac, got it?

OpenStudy (anonymous):

yes so first we divid everything by 3?

OpenStudy (anonymous):

if you do that, you'll get b/3

OpenStudy (anonymous):

yes is that correct?

OpenStudy (anonymous):

hmm...show me how you do

OpenStudy (anonymous):

i dont get it :'(

OpenStudy (anonymous):

if i divid by 3 ill get x^2+bx+9 i take c multiply it by 2 :D

OpenStudy (anonymous):

you'll get x^2+(b/3)x+9

OpenStudy (anonymous):

i dont get your way its so complicated :/ where did you get tryx from

OpenStudy (anonymous):

random constants xD

OpenStudy (anonymous):

lol choose x lol